OpenStudy (anonymous):

limit as x approaches 1+ ((1/x-1)-(1/lnx))

myininaya (myininaya):

\[\lim_{x \rightarrow 1^+} (\frac{1}{x}-1-\frac{1}{\ln(x)}) \]

OpenStudy (anonymous):

i think it is \[\lim_{x\rightarrow 1^+}\frac{1}{x-1}-\frac{1}{\ln(x)}\]

OpenStudy (anonymous):

but i could be wrong for sure

myininaya (myininaya):

it is up to interpretation i believe its like art

OpenStudy (anonymous):

i know it is equal to -1/2, but I'm not sure of the steps.

myininaya (myininaya):

would you like to tell us what the problem is? does satellite have it right?

OpenStudy (anonymous):

it would help if we knew it was what i wrote or what myininaya wrote, because they are different

OpenStudy (anonymous):

sorry! what satellite wrote. damn typos...

OpenStudy (anonymous):

i get \[-\frac{1}{2}\]

OpenStudy (anonymous):

start with \[\frac{\ln(x)-x+1}{x\ln(x)-\ln(x)}\] then use l'hopital twice

myininaya (myininaya):

\[\lim_{x \rightarrow 1^+}\frac{\ln(x)-(x-1)}{\ln(x) \cdot (x-1)}=\lim_{x \rightarrow 1^+}\frac{\ln(x)-x+1}{x \ln(x)-\ln(x)}\] since if we plug in 1 we get 0/0 we can use lhospital \[=\lim_{x \rightarrow 1}\frac{\frac{1}{x}-1}{\ln(x)+x \cdot \frac{1}{x}-\frac{1}{x}}\] \[=\lim_{x \rightarrow 1}\frac{\frac{1}{x}-1}{\ln(x)-\frac{1}{x}+1}\] we can do it again since we have 0/0 again \[=\lim_{x \rightarrow 1}\frac{\frac{-1}{x^2}}{\frac{1}{x}-\frac{-1}{x^2}} =\lim_{x \rightarrow 1}\frac{-1}{x+1}\]

myininaya (myininaya):

=-1/2

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

thanks so so much.

OpenStudy (anonymous):

at step two feel free to multiply top and bottom by x to turn \[=\lim_{x \rightarrow 1}\frac{\frac{1}{x}-1}{\ln(x)-\frac{1}{x}+1}\] into \[\frac{1-x}{x\ln(x)+x-1}\] then use l'hopital again

myininaya (myininaya):

why i don't have to use product rule with mine after that one step

OpenStudy (anonymous):

myininaya usually loves to clear fractions and i do not. we traded places this time

myininaya (myininaya):

i like to avoid product rule when i can and just do simple derivatives where power rule is only thing involved

OpenStudy (anonymous):

yeah but you already found the derivative of \[x\ln(x)\] in the line above so you get it for free! it is not like it changed from one line to the next

myininaya (myininaya):

lol

myininaya (myininaya):

it requires more writing though so it does cost

OpenStudy (anonymous):

and no one uses the power rule to find the derivative of \[\frac{1}{x}\] it is like using a calculator to multiply by 5

myininaya (myininaya):

i wish zarkon was here so he could agree with me

OpenStudy (anonymous):

you could always just delete my post

myininaya (myininaya):

my way is easiest! i win!

OpenStudy (anonymous):

my way is easiest.... you win anyway

myininaya (myininaya):

good game

OpenStudy (zarkon):

you two are too funny

myininaya (myininaya):

who wins zarkon?

OpenStudy (anonymous):

i do of course

myininaya (myininaya):

no!

OpenStudy (anonymous):

(the arrogance!)

OpenStudy (zarkon):

a solid tie

myininaya (myininaya):

no way!

OpenStudy (zarkon):

that's what the score card says. I only report what the cards say

OpenStudy (anonymous):

rubber match tomorrow

myininaya (myininaya):

it doesn't matter to you which one: (xln(x))' or (1/x)' :( power rule is easiest!

myininaya (myininaya):

product rule requires more writing

OpenStudy (zarkon):

if you know the taylor expansion at x=1 for the ln(x) the solution it pretty short

myininaya (myininaya):

lol

myininaya (myininaya):

goodnight zarkon